A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.
Each employee has 8 choices of drink,
so total # of ways they can choose = 8^6
ways in which all take different drinks = 8P6 = 876..3
so ways in which at least 2 take the same drink = 8^6 - 8P6
probability = (8^6 - 8P6) /8^6
= .9231
I hope my answer has come to your help. Thank you for posting your question here in Brainly. We hope to answer more of your questions and inquiries soon. Have a nice day ahead!
Answer:
0.923
Step-by-step explanation:
Given:
=> the total possible outcomes: \(8^{6}\)
The probability that at least 2 employees purchased the same drink, it means that we can use: 1 - The probability that at no employees purchased the same drink.
So, number favorable outcomes at no employees purchased the same drink: 876543
Hence, The probability that at no employees purchased the same drink:
= number favorable outcomes / the total possible outcomes
= 876543 / \(8^{6}\)
= 0.77
=> The probability that at least 2 employees purchased the same drink: 1- 0.77 = 0.923